Flint Co. invested $970,000 in Buffalo Co. for 25% of its outstanding stock. Buffalo Co. pays out 40% of net income in dividends each year. Use the information in the following T-account for the investment in Buffalo to answer the following questions.

Investment in Buffalo Co.

970,000    48,000

120,000

(a) How much was Flint Co.’s share of Buffalo Co.’s net income for the year? Net income $enter Flint’s net income in dollars 120000

(b) What was Buffalo Co.’s total net income for the year? Total net income $enter the total net income of BuffaloCo. for the year 48000

(c) What was Buffalo Co.’s total dividends for the year? Total Dividends $enter the total dividends of BuffaloCo. for the year 480000

(d) How much was Flint Co.’s share of Buffalo Co.’s dividends for the year? Dividends $enter the dollar amount of dividends for Flint Co.
Answer #1

a) Flint Co's share in Buffalo Co = 120000

b) Total net income = 120000*100/25 = 480000

c) Total dividend = 480000*40% = 192000

d) Flint's Share = 192000*25% = 48000